A metal sulfide with the empirical formula of MS 2 , contains 40.06 % sulfur

by mass. Calculate the molar mass of M.


1
Expert's answer
2021-09-03T01:47:32-0400

Solution:

Molar mass of S = 32.065

Suppose the molar mass of M is x, then:

Molar mass of MS2 = x + 2×32.065 = x + 64.13


%S / 100% = (2 × Molar mass of S) / (Molar mass of MS2)

40.06% / 100% = (2 × 32.065) / (x + 64.13)

0.4006 = (64.13) / (x + 64.13)

160.08 = x + 64.13

Molar mass of M = x = 160.08 - 64.13 = 95.95

Molar mass of M = 95.95

Hence, M = Molybdenum (Mo)

MS2 = MoS2


Answer: The molar mass of M is 95.95 g/mol